Example document

markus schnalke

2014-05-05

1 Text

The Debian GNU/Linux distribution [1] is completely community-driven.

The definitive introduction to the C programming language is Kernighan and Ritchie’s book The C Programming Language [KR88].

An overview on the history of the Internet can be found in the Wikipedia [2].

Literatur

[KR88] Brian W. Kernighan and Dennis M. Ritchie. The C programming language. Prentice-Hall, second edition, 1988. ISBN: 0-13-110362-8.

Websites

[1] The Debian Project. Debian website. http://debian.org (2009-01-15).

[2] Wikipedia. History of the internet. http://en.wikipedia.org/w/index.php?title=
   History_of_the_Internet&oldid=245178555 (2008-10-23).